CENGIZ ERBAS

Work Address TRW, 300 Lakeside Drive, Oakland, California 94612

(510) 645–3556 Home Address 3156 Oak Road, #201 Walnut Creek, California 94596 (510) 210–0876

Education  Southern Methodist University  Dallas, Texas

Ph.D. in Computer Science, December 1994, (GPA = 3.94). Dissertation: Modeling and Verification of Real-time Systems.

Southern Methodist University Dallas, Texas M.S. in Computer Science, August 1992, (GPA = 3.96). Thesis: N-Queens Problem and Its Applications in Computer Science.

Hacettepe University Ankara, Turkey B.S. in Computer Science, June 1989, (GPA = 3.51, Summa Cum Laude).

Experience  TRW Financial Systems  Oakland, California

April 1994–Present Responsible team lead for the design and development of the VME-based real-time transport control unit (TRCU) of TRW image processing systems for a document recognition project. Developed a string pattern matching algorithm (for real-time parsing of OCRlines) with an upper bound O((f_{i} + f_{j})n), where f_{i} and f_{j} are two flexibility constants for the insert and delete operations, and n is the length of the OCR string. Integrated the parser with the rest of the image system, and demonstrated that the algorithm satisfies the timing constraints of TRCU.

Superconducting Super Collider Laboratory Dallas, Texas January 1992–April 1994 Participated in the development of the real-time control systems of SSCL Accelerators. Worked on Input/Output Controller of EPICS. Developed several device drivers on the VxWorks real-time kernel for various VME boards.

Participated in the development and integration of a UNIX-based portable data acquisition system (UNIDAQ) written in C and C++. Developed the underlying generic template structure which handles the communication and synchronization of system processes, and integrated the template with the other components of the system. Ported UNIDAQ from UNIX to VxWorks. Ported a CAMAC device driver K2917-K3922 from SunOS 4.1.3 to MVME-147 and MVME-167 modules, which run the VxWorks real-time kernel.

Systems Software Engineering Laboratory, SMU Dallas, Texas August 1990–January 1992 Participated in research on expert systems, parallel memory systems, and computer applications of the n-queens problem.

Software/Hardware Background Experience on UNIX, VxWorks, Solaris 2.x, POSIX, AOS/VS, VAX/VMS, BTOS, DOS.

C, C++, Pascal, FORTRAN 77, Smalltalk/V, COBOL, INTEL 8086/8088, INTEL 8080A/8085, PROLOG, BASIC.

Experience on TCP/IP, X-Windows, PAW, EPICS, DataViews, GKS, SQL, SYBASE.

SUN Workstations, M68040, M68030, VMEbus, VAX 11/780, Data General MV/10000, UNISYS B27, UNISYS XE-520, IBM PC’s and Macintosh’s.

Journal Publications “Parallel Memory Allocation and Data Alignment in SIMD Machines,” Journal of Parallel Algorithms and Applications, 4(1994), 139-151. [with M.M. Tanik and V.S.S. Nair]

“Linear Congruence Equations for the Solutions of the N-Queens Problem,” Information Processing Letters, 41(1992), 301-306. [with M.M. Tanik and Z. Aliyazicioglu]

“Selection of Expert System Development Tools for Engineering Applications,” Transactions of the ASME, Journal of Energy Resources Technology. (March 1992). [with S. Mills, Y. Hurmuzlu and M.M. Tanik]

Conference Publications “A Precedence Graph Model for Real-Time Systems,” IEEE COMPSAC 94, Computer Software and Applications Conference, Taipei, Taiwan, (November 9-11, 1994). [with M.M. Tanik]

“A Theoretical Foundation for Rapid Prototyping of Time Critical Systems,” ASME 1994 ESDA Conference, London, England, (July 4-7, 1994). [with M.M. Tanik, A. Ertas and S.G. Gatchel]

“A Circulant Matrix based approach to Storage Schemes for Parallel Memory Systems,” 5th IEEE Symposium on Parallel and Distributed Processing, Dallas, Texas, (December 1-4, 1993). [with M.M. Tanik and V.S.S. Nair]

“A Knowledge Based Advisor for the Selection of Expert System Development Tools,” ASME 1992 ESDA Conference, Istanbul, Turkey, (June 29-July 3, 1992). [with S. Mills and M.M. Tanik]

“Different Perspectives of the N-Queens Problem,” ACM 1992 Computer Science Conference, Kansas City, MO, (March 3-5, 1992). [with S. Sarkeshik and M.M. Tanik]

“Storage Schemes for Parallel Memory Systems and the N-Queens Problem,” 15th Annual ASME ETCE Conference, Computer Applications Symposium, Houston, Texas, (January 26-30, 1992). [with M.M. Tanik]

“Engineering Problem Specification with Smalltalk/V,” 14th ASME ETCE Conference, Workshop on Expert Systems and Applications, Houston, Texas, (January 20-23, 1991). [with E. Demirors, O. Demirors, C. Bayrak and M.M. Tanik]

Technical Reports “Observations on Modeling and Verification of Real-Time Systems,” SMU Technical Report 93-CSE-53, (November 1993). [with M.M. Tanik and A.C. Fry]

“Writing Device Drivers for the VxWorks Operating System, A Case Study: MVME167-CAMAC and MVME147-CAMAC Device Drivers,” Superconducting Super Collider Laboratory, (May 1992). [with M. Botlo and A. Fry]

“N-Queens Problem and Its Connections to Polygons,” SMU Technical Report 91-CSE-21, June 1991. [with M.M. Tanik]

“N-Queens Problem and Its Algorithms,” SMU Technical Report 91-CSE-8, (February 1991). [with M.M. Tanik] Presentations “SDC Portable Data Acquisition System: UNIDAQ,” KEK (National Laboratory for High-

Energy Physics), Tsukuba, Japan, (March 2, 1993).

“On-line Computing Requirements of Superconducting Super Collider Laboratory,” High Performance Computing Center, University of Texas at Austin, (August 10, 1992).

“N-Queens Problem, Its Algorithms and Some Observations on the Backtracking Search Tree of the N-Queens Problem,” 15th Computer Science Conference, Hewlett-Packard, Denton, TX, (April 26, 1991).

Awards “TRW Recognition Award,” of TRW Financial Systems, Oakland, California, 2 times: (October 1994) and (January 1995).

“1994 Science Award,” of Informatics Association of Turkey, Istanbul, Turkey, (September 1994).

“Outstanding Graduate Student Award,” of Computer Science and Engineering Department of Southern Methodist University, Dallas, Texas, (April 1994).

“Donald Edwards PhD Dissertation Scholarship,” of Superconducting Super Collider Laboratory, Dallas, Texas, (January 1993-May 1994).

“Outstanding Achievement Award,” of Dogramaci Foundation, Ankara, Turkey, (September 1989).

“Scientific Encouragement Award,” of Hacettepe University, Ankara, Turkey, (September 1989).

Service and Extracurricular Activities Session Co-chairman of 17th ASME Energy-sources Technology Conference, Computers in Engineering Symposium, Computer Applications Session, New Orleans, Louisiana, (January 24, 1994), ASME, PD-Vol. 59.

Session Co-chairman of 17th ASME Energy-sources Technology Conference, Computers in Engineering Symposium, Computer-Aided Modeling, Design and Manufacturing Session, New Orleans, Louisiana, (January 26, 1994), ASME, PD-Vol. 59.

President of the ACM Student Chapter at SMU, Dallas, TX, (January 1991- January 1992).

Member of ACM, IEEE, ASME, Planetary Society, Informatics Association of Turkey.